Newton's Second Law of Motion


Show your solution.


A truck hauls a car cross-country. The truck's mass is 4.00x10 ³kg and the car's mass is 1.60x10 ³kg. if the force of propulsion resulting from the truck's turning wheels is 2.50x10à ¢  ´N, then determine the acceleration of the car (or the truck) and the force at which the truck pulls upon the car. Assume negligible air resistance forces.


1
Expert's answer
2021-10-20T10:40:51-0400
"a=\frac FM=\frac{2.50·10^3}{4·10^3+1.6·10^3}=0.45\text{ m/s}^2."